OpenStudy (anonymous):

Which point lies on a circle with a radius of 5 units and center at P(6, 1)? Q(1, 11) R(2, 4) S(4, -4) T(9, -2)

OpenStudy (anonymous):

the equation for a circle is r^2=x^2+y^2 So your radius is 5 and you have x and y values (both positive, so we will subtract them) which gives us the equation 25=(x-1)^2 + (y-6)^2 Now if you took the possible answers and plugged them into the equation you can see which point is correct, because the correct x and y values would equal 25. try it out and tell me what you think the answer might be.

OpenStudy (anonymous):

Also this should be posted in geometry section